OpenStudy (anonymous):

Find the surface area of the ball. Leave the answer in terms of pie. d = 24 cm

OpenStudy (anonymous):

d = 2*r, so r = d/2 = 24/2 = 12cm

OpenStudy (anonymous):

Volume of a ball is V = (4/3)pi*r^3 = (4/3)pi*144*12 = 16*144pi = 2304pi

OpenStudy (anonymous):

surface area of a ball is 4 x pie x r^2

OpenStudy (anonymous):

Oh, sorry I misread it. I thought they asked for the volume..

OpenStudy (anonymous):

SA = 4*pi*r^2 = 4*144*pi=576pi cm^2

OpenStudy (anonymous):

yeah, unusual to be asked surface area really in this kinda thing

OpenStudy (anonymous):

thanks!